# AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

### v1

@@ Appendix H Empirical Analysis of the Approximate Submodularity of Function $F$
To offer a more comprehensive insight into the motivation for employing a greedy algorithm, this section provides a thorough analysis. We emphasize that for functions that are not theoretically submodular, a greedy approach remains one of the conventional methods for addressing their maximization, as established in recognized works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011). In this part, we conduct meticulously designed experiments to investigate the extent to which the information gain function $F$ approximates submodularity. The experimental results reveal that $F$ empirically exhibits significant submodular characteristics. This finding motivates us to follow established works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011) to propose a greedy algorithm to solve the maximization problem for $F$. The analysis is detailed as follows:
@@
for any sets $A\subseteq B\subset C$ and any element that satisfies $R_{i}\in C\backslash B$. In our scenario, the Inequality 15 is written as
@@
Next, to compute the information gain contributed by a given region, we randomly sample a region $R_{test}$ from $C\setminus S_{large}$. We then calculate the VLM’s information content, which are denoted as $U_{s},U_{s}^{*},U_{l},\text{and }U_{l}^{*}$, when the context incorporates (1) $S_{small}$, (2) $S_{small}\cup\{R_{test}\}$, (3) $S_{large}$, and (4) $S_{large}\cup\{R_{test}\}$, respectively. We expect to observe in the majority of cases that:
@@
Experimental Results. The experimental results are presented in Table 6. As we can see, the Inequality 15 holds in most instances across all settings and datasets. Furthermore, to confirm the significance of the obtained results, we introduce the Binomial Test, an exact statistical procedure for assessing the extent to which experimental outcomes with a binary structure are attributable to chance alone. The p-values, presented in Table 6, are all substantially below the 0.05 significance level. This demonstrates that the information gain function $F$ behaves in a manner that is empirically near-submodular, which motivates us to follow existing research Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011) where greedy algorithms are proposed to solve the problem of maximizing approximately submodular functions.
Table 6: Proportions of instances on M3CoT and LLaVA-W benchmarks for which the approximate submodularity of information gain function $F$ is manifested. The backbone is Chameleon-7B and the model is our proposed AIMCoT. $K_{large}$ is set to $K_{small}+1$ for simplicity. The significance levels of these results are listed below them.
@@
Appendix I Analysis of the Selected Regions’ Source
In this section, we examine the distribution of sources for the visual regions selected by the AVP module of AIMCoT. These regions are drawn from two sets, $C_{attn}$ and $C_{exp}$, with their respective selection proportions denoted as $P_{attn}$ and $P_{exp}$. Intuitively, $P_{exp}$ reflects the significance of incorporating the exploratory set $C_{exp}$ to construct a better multimodal CoT. A larger value of $P_{exp}$ indicates that the exploratory set $C_{exp}$ makes a greater contribution by providing informative salient regions to AIMCoT, and vice versa.
Experimental Setup The experiments are conducted on the M3CoT and LLaVA-W benchmarks. Our proposed AIMCoT is implemented with the Chameleon-7B backbone under a default 0-shot setting. To ensure the reliability of the results, we repeat each experiment three times on both benchmarks.
Results and Analysis As presented in Table 7, although the value of $P_{exp}$ fluctuates across different experimental runs on the same benchmark, it remains consistently around 20% on M3CoT and 30% on LLaVA-W. This indicates that the influence of stochastic factors on the source distribution of the selected regions is limited, which validates our rationale of using this metric as a reflection of the relative importance of $C_{attn}$ and $C_{exp}$. Furthermore, we observe that $P_{exp}$ is significantly greater than zero. This demonstrates that the exploratory set $C_{exp}$ consistently serves as a critical component of the total candidate set $C$, contributing a substantial portion of the informative regions for AIMCoT.
Table 7: Proportion of salient regions selected by the AVP module of our proposed AIMCoT from the exploratory set $C_{exp}$.
Experiment Number 1 2 3
M3CoT 17.25% 20.44% 27.27%
LLaVA-W 31.33% 25.77% 26.67%
Appendix J Deployment of AIMCoT
J.1 Analysis of the Complexity of AVP Module
@@ Appendix I Analysis of the Selected Regions’ Source
The Active Visual Probing (AVP) module’s primary purpose is to dynamically and intelligently select salient sub-regions of an image during the text generation process. This is achieved by calculating the “information gain” that each potential sub-region offers, thereby allowing the model to “zoom in” on relevant visual details and generate more informed and contextually aware text.
@@
This formula highlights that, thanks to KV caching, the complexity scales linearly with sequence length N—a substantial improvement over the standard quadratic dependence. While the cost is also linear with respect to the number of candidates $N_{C}$ and selections $K$, our framework maintains strong deployability. This is because the AVP module is highly efficient at extracting crucial visual information from the candidate set. Our empirical results demonstrate that AIMCoT achieves exceptional performance (as shown in Table 2) even when these hyper-parameters are kept at low levels (e.g., $K=3,N_{C}=6$, which is the default setting). In conclusion, we can approximate the complexity of AVP as follows:
@@
This demonstrates that the combination of architectural optimizations (batching and KV caching) and the high extractive efficiency of AVP ensures the module’s practicality, making it efficient and readily deployable in practice. As a comparison, the attention-driven selection method within the baseline model ICoT is with a complexity of $O(logK\cdot\Delta N+L\cdot N\cdot H)$, which also scales linearly with sequence length N. Although this is acknowledged to be lower than the complexity of AVP shown in Equation 19, our empirical results detailed in Appendix J.2 suggest that AVP’s average inference time is no more than 1.36 times that of this method, while achieving performance far superior to it as analyzed in Section 5.5.
J.2 Empirical Analysis of the Deployment of AIMCoT
@@
In this section, we empirically investigate the deployability of the AIMCoT framework and the temporal overhead introduced by the AVP module. For the experimental setup, we utilized Chameleon-7B as the backbone in a 0-shot setting to compare the average inference time of AIMCoT against ICoT Gao et al. (2025). ICoT, as a key baseline model in this study, employs a Top-K strategy to simultaneously select regions with the highest attention scores for constructing the multimodal CoT, thereby expected to possess a relatively lower time complexity compared to AIMCoT. Consequently, the comparison with ICoT serves as a direct indicator of AIMCoT’s deployability.
Table 8: Comparison of the average time to process each instance between AIMCoT and the baseline model (ICoT).
Dataset AIMCoT ICoT
M3CoT 13.37s 11.62s
LLaVA-W 11.65s 8.58s
@@
The experimental results are presented in Table 8. Two key observations can be drawn. First, the AVP module does not introduce significant temporal costs to the AIMCoT framework, an efficiency attributable to batch processing and the KV Cache mechanism. Second, AIMCoT achieves substantially superior performance as shown in Table 2, at a time cost comparable to that of the efficient baseline, which is less than $1.36$ times that of ICoT (specifically, $1.15$ and $1.36$ times on M3CoT and LLaVA-W benchmarks, respectively). This suggests that our proposed AIMCoT framework achieves a favorable trade-off between performance and deployability.
Appendix K Sensitivity Analysis of $\delta$
The hyper-parameter $\delta$ within the DAT module serves as a crucial threshold to trigger the AVP module, which inserts salient visual regions to improve the construction of the multimodal CoT. In this section, we detail a sensitivity analysis of $\delta$ by adjusting $\delta$ across the range of $[0.1,0.125,0.15,0.175,0.2,0.225]$ and examining not only (1) the performance of AIMCoT, but also (2) the number of times the AVP is triggered. The experiments are conducted under 0-shot setting on Chameleon-7B backbone and LLaVA-W benchmark. The experimental results are shown in Figure 5.
The left figure illustrates that AIMCoT exhibits limited performance when the threshold, $\delta$, is set too low. This underscores the importance of inserting visual information at critical moments: excessively frequent or inopportune visual insertions can disrupt the VLM’s reasoning process, leading to suboptimal performance. As $\delta$ increases, the model’s performance progressively improves, reaching its peak at $\delta=0.2$ (our default setting), which corresponds to a ROUGE-L score of 0.2983. However, a further increase in $\delta$ results in a slight performance degradation. This highlights the criticality of visual information insertion for constructing an interleaved Chain of Thought: an overly stringent threshold excessively impedes the incorporation of visual data, preventing the AVP from supplying the model with necessary visual supplementation in a timely manner.
Figure 5: Experimental results of the sensitivity analysis of the hyper-parameter $\delta$. The left figure illustrates the performance of AIMCoT when $\delta$ takes different values, while the right one shows the number of times the AVP module within AIMCoT is triggered.
Table 9: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on M3CoT benchmark, which illustrate the average time of AIMCoT to process an instance under different settings. Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
$N_{C}$ 1 2 3 4 5 6 7
K=1 12.78s 12.80s 12.83s 12.87s 12.90s 12.93s 12.97s
K=3 - - 13.18s 13.23s 13.30s 13.37s 13.42s
K=5 - - - - 13.41s 13.49s 13.56s
K=7 - - - - - - 13.66s
Table 10: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on LLaVA-W benchmark, which illustrate the average time of AIMCoT to process an instance under different settings. Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
$N_{C}$ 1 2 3 4 5 6 7
K=1 10.61s 10.70s 10.75s 10.84s 10.92s 11.00s 11.08s
K=3 - - 11.21s 11.36s 11.51s 11.65s 11.80s
K=5 - - - - 11.79s 11.96s 12.14s
K=7 - - - - - - 12.38s
Appendix L Sensitivity Analysis of $K,N_{C}$
Our proposed AIMCoT incorporates the design of AVP module. In contrast to existing research, AIMCoT, benefiting from the AVP module, does not simply select the top-$K$ regions with the highest attention scores from the attention map. Instead, it meticulously selects $K$ regions from a total set $C$ of $N_{C}$ candidate regions to construct the multimodal CoT ($N_{C}=N+M$). However, this approach may inevitably raise concerns regarding the deployability of AIMCoT, particularly as the hyper-parameters $N_{C}$ and $K$ increase.
To investigate this, we conduct experiments to explore the average processing time per instance for AIMCoT with larger values of $N_{C}$ and $K$. For the experimental setup, we implement AIMCoT with the Chameleon-7B backbone on the M3CoT and LLaVA-W datasets under various combinations of $N_{C}$ and $K$. The results are presented in Tables 9 and 10, from which we derive two key insights:
  • Insensitivity to the growth of the candidate set size $N_{C}$: observing any row with a fixed $K$ (e.g., $K=3$ on Table 9), as $N_{C}$ increases from 3 to 7, the total processing time rises from 13.18s to 13.42s, a marginal increase of only 0.24s. This implies that each additional candidate region introduces an average overhead of less than 0.06s. This strongly demonstrates that the performance of the AVP module does not degrade sharply with a moderate expansion of the candidate pool, indicating excellent scalability.
  • Diminishing marginal cost with the increase in the number of selections $K$: considering a fixed column for $N_{C}$ (e.g., $N_{C}=7$ on Table 10), as $K$ increases from $1\rightarrow 3$, $3\rightarrow 5$, and $5\rightarrow 7$, the processing time increases by 0.72s, 0.34s, and 0.24s, respectively, which shows a notable decrease in incremental cost. This suggests that the primary computational overhead of the AVP algorithm lies in the initiation of the iterative search (the jump from $K=1$ to $K=3$). Once the iteration begins, the cost of subsequent selection steps is remarkably low, benefiting from efficient mechanisms such as the batch processing and KV Cache.
Table 11: Performance comparison AIMCoT variants on the basis of different constructions of the candidate set $C$.
Construction of $C$ M3CoT (ACC.) LLaVA-W (ROUGE-L)
$C_{attn}\cup C_{rand}~~~(C_{exp}=C_{rand})$ 31.4 29.8
$C_{attn}\cup C_{ss}~~~(C_{exp}=C_{ss})$ 31.2 29.5
$C_{attn}\cup C_{fsam}~~~(C_{exp}=C_{fsam})$ 31.0 29.6
$C_{attn}$ 30.8 28.9
$C_{rand}$ 30.4 28.6
$C_{ss}$ 30.3 28.7
$C_{fsam}$ 29.9 27.7
Appendix M Ablation Study on the Construction of Candidate Set $C$
In this section, we investigate the influence of different compositions of the total set $C$ on the performance of our proposed AIMCoT. We specifically examine two primary configurations:
Constructing $C$ using only $C_{attn}$ or $C_{exp}$. For the latter, we evaluate four distinct construction methodologies for $C_{exp}$: (a) $C_{rand}$: uniform random sampling; (b) $C_{ss}$: the selective search algorithm Uijlings et al. (2013), which is the seminal region proposal method utilized in R-CNN Girshick et al. (2014); (c) $C_{fsam}$: FastSAM Zhao et al. (2023), a computationally efficient variant of the foundational vision segmentation model, SAM Kirillov et al. (2023).
As observed, the combination of the two sets (i.e., $C=C_{attn}\cup C_{exp}$) invariably yields superior performance for AIMCoT compared to configurations where either $C=C_{attn}$ or $C=C_{exp}$ is used exclusively. This highlights the importance of diversifying the sources of candidate visual regions.

### v2

@@ I.4 Theoretical Justification: Approximate Submodularity and Optimality of Greedy Selection
To offer a more comprehensive insight into the motivation for employing a greedy algorithm, this section provides a thorough analysis. We emphasize that for functions that are not theoretically submodular, a greedy approach remains one of the conventional methods for addressing their maximization, as established in recognized works Bian et al. (2017); Sener and Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das and Kempe (2011). In this part, we conduct meticulously designed experiments to investigate the extent to which the information gain function $F$ approximates submodularity. The experimental results reveal that $F$ empirically exhibits significant submodular characteristics. This finding motivates us to follow established works Bian et al. (2017); Sener and Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das and Kempe (2011) to propose a greedy algorithm to solve the maximization problem for $F$. The analysis is detailed as follows:
@@
for any sets $A\subseteq B\subset C$ and any element that satisfies $R_{i}\in C\backslash B$. In our scenario, the Inequality 12 is written as
@@
Next, to compute the information gain contributed by a given region, we randomly sample a region $R_{test}$ from $C\setminus S_{large}$. We then calculate the VLM’s information contents, which are denoted as $U_{s},U_{s}^{*},U_{l},\text{and }U_{l}^{*}$, when the context incorporates (1) $S_{small}$, (2) $S_{small}\cup\{R_{test}\}$, (3) $S_{large}$, and (4) $S_{large}\cup\{R_{test}\}$, respectively. We expect to observe in the majority of cases that:
@@
Experimental Results. The experimental results are presented in Table 16. As we can see, the Inequality 12 holds in most instances across all settings and datasets. Furthermore, to confirm the significance of the obtained results, we introduce the Binomial Test, an exact statistical procedure for assessing the extent to which experimental outcomes with a binary structure are attributable to chance alone. The p-values, presented in Table 16, are all substantially below the 0.05 significance level. This demonstrates that the information gain function $F$ behaves in a manner that is empirically near-submodular, which motivates us to follow existing research Bian et al. (2017); Sener and Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das and Kempe (2011) where greedy algorithms are proposed to solve the problem of maximizing approximately submodular functions.
Table 16: Proportions of instances on M3CoT and LLaVA-W benchmarks for which the approximate submodularity of information gain function $F$ is manifested. The backbone is Chameleon-7B and the model is our proposed AIM-CoT. $K_{large}$ is set to $K_{small}+1$ for simplicity. The significance levels of these results are listed below them.
@@
Appendix J Extensive Ablation Studies on More Backbones
Section 5.3 presents an ablation study on the Chameleon-7B backbone. The results provide two key insights as follows:
  • Each component (CAG, AVP, DAT) contributes substantially to the overall performance.
  • Even without CAG, combining AVP and DAT is sufficient for AIM-CoT to outperform strong baseline models by a clear margin. This finding highlights the importance of addressing the “what to see” and “when to see it” questions.
To test whether these insights generalize beyond Chameleon-7B, in this section, we extend the same ablation study to two additional VLM backbones: Janus-Pro-7B and Qwen2.5-VL-32B.
As shown in Tables 17 and 18, the ablation trends are consistent with those in Section 5.3: removing any component leads to performance degradation across datasets and backbones, demonstrating the efficacy of our proposed components.
More importantly, the results highlight the central thesis of this study. Even without CAG, the combination of AVP and DAT already yields strong performance and is sufficient to surpass strong baselines, including ICoT. This demonstrates that explicitly and effectively addressing what to see (AVP) and when to see it (DAT) is the key driver of AIM-CoT’s advantage.
Meanwhile, CAG remains a non-trivial contributor: although its absolute gains are comparatively smaller, it consistently improves results across all settings, suggesting that mitigating the text-vision granularity mismatch further complements AVP+DAT.
Table 17: Ablation study of AIM-CoT conducted on Janus-Pro-7B under 0-shot setting.
Dataset AIM-CoT w/o CAG w/o AVP w/o DAT
M3CoT (ACC.) 39.7 39.4 (-0.3) 38.6 (-1.1) 39.0 (-0.7)
ScienceQA (ACC.) 56.9 56.7 (-0.2) 55.8 (-1.1) 56.3 (-0.6)
LLaVA-W (ROUGE-L) 35.5 34.7 (-0.8) 33.5 (-2.0) 34.5 (-1.0)
Table 18: Ablation study of AIM-CoT conducted on Qwen2.5-VL-32B under 0-shot setting.
Dataset AIM-CoT w/o CAG w/o AVP w/o DAT
M3CoT (ACC.) 58.7 58.4 (-0.3) 57.6 (-1.1) 58.1 (-0.6)
ScienceQA (ACC.) 76.8 76.6 (-0.2) 75.9 (-0.9) 76.3 (-0.5)
LLaVA-W (ROUGE-L) 46.5 46.1 (-0.4) 44.2 (-2.3) 45.0 (-1.5)
Appendix K In-depth Analysis of Performance Gain Variations Across Backbones
In Table 1, we observe that AIM-CoT yields varying degrees of improvement across different base models. Specifically, the gains on Chameleon-7B are substantially larger than those on Qwen2-VL-7B, while Qwen2.5-VL-32B exhibits higher improvements than its 7B counterpart. In this section, we analyze the underlying factors contributing to these phenomena from two perspectives: visual encoding mechanisms and model capability scaling.
K.1 Cross-Family Analysis: Architecture and Visual Encoding
The distinct performance gap between Chameleon-7B (up to 18.25% gain on LLaVA-W) and Qwen2-VL-7B (up to 6.14% gain) can be attributed to their fundamental differences in visual processing:
  • Chameleon-7B (Early-Fusion & Fixed Resolution): As an early-fusion model utilizing a fixed tokenization strategy, Chameleon-7B often struggles with fine-grained visual details in high-resolution images, leading to the “myopic” attention behavior discussed in Section 3.1. AIM-CoT acts as a perceptual correction mechanism here. By actively cropping and zooming in on salient regions via the AVP module, AIM-CoT directly compensates for the model’s native resolution limitations. The improvement is transformative as it solves a “can’t see” problem, resulting in substantial gains.
  • Qwen2-VL Series (NaViT & Dynamic Resolution): These models employ the NaViT architecture with dynamic resolution support, allowing them to process arbitrary aspect ratios and resolutions natively. Since Qwen2-VL can already perceive visual details relatively clearly, the marginal utility of “zooming in” is lower compared to Chameleon. For Qwen models, AIM-CoT serves more as an attention optimization tool rather than a vision repair tool, leading to more moderate base improvements.
K.2 Intra-Family Analysis: Capability Dependence and Scaling Laws
Within the Qwen family, we observe that the larger Qwen2.5-VL-32B benefits more from AIM-CoT (up to 9.84% gain) than the smaller Qwen2-VL-7B (up to 6.14% gain). This counter-intuitive trend (where a stronger baseline yields larger relative gains) highlights the capability-dependent nature of our framework:
  • Quality of CAG Generation: AIM-CoT is a closed-loop system where the Context-enhanced Attention-map Generation (CAG) module relies on the model’s own capability to describe the image. The 32B model generates more precise, hallucination-free descriptions than the 7B model. This high-quality context leads to more accurate attention maps, enabling the AVP module to select significantly more valuable image regions.
  • Interleaved Reasoning Capability: The efficacy of AIM-CoT also depends on how well the model can leverage the inserted visual patches. Larger models (32B) possess stronger long-context understanding and multi-step reasoning capabilities. They can effectively synthesize the fragmented information provided by the inserted patches to deduce the correct answer, whereas smaller models might struggle to integrate these additional visual cues coherently.
In summary, AIM-CoT functions as a mechanism for Perceptual Correction on weaker architectures (Chameleon) and Reasoning Augmentation on stronger ones (Qwen 32B), demonstrating its adaptability across different model paradigms.
Appendix L Safety Mechanisms in AIM-CoT
To ensure the reliability of the reasoning chain and prevent error propagation, AIM-CoT incorporates explicit safety mechanisms at two critical stages. First, regarding Source Quality (Input), we apply Negative Constraints within the CAG module to strictly ground the generated context and mitigate hallucinations. Second, regarding Integration Safety (Process), we employ Safety Instructions during the visual insertion, guiding the model to verify relevance and filter out potential noise. In this section, we analyze the individual contributions of these safeguards to the framework’s overall robustness.
L.1 Effectiveness of Negative Constraints in CAG Generation
In Section 4.2, we introduce negative constraints within the CAG prompt $\mathcal{P}_{CAG}$ to ensure the utmost precision of the generated descriptions. To rigorously assess the necessity of this design and the potential risk of error propagation, we conduct an ablation study focusing on the hallucination rate of the generated descriptions $\mathcal{D}_{CAG}$ and their impact on downstream performance.
Experimental Setup. We contrast the standard AIM-CoT against AIM-CoT w/o Constraints, a variant where the cautionary instructions are removed from $\mathcal{P}_{CAG}$. To demonstrate the generalizability of our design, we implement the framework on two representative backbones: Chameleon-7B and Qwen2-VL-7B. The experiments are conducted on M3CoT, ScienceQA, and LLaVA-W benchmarks under the 0-shot setting.
Evaluation Metrics. Two metrics are incorporated for evaluation: (1) Hallucination Rate (HR): We employ GPT-4v as an external evaluator to identify factual inconsistencies between the image content and the generated description $\mathcal{D}_{CAG}$. HR is reported as the percentage of descriptions containing details unsupported by the visual input. (2) Downstream Performance: We report Accuracy (ACC.) for M3CoT and ScienceQA, and ROUGE-L for LLaVA-W to measure the final reasoning outcome.
Table 19: Ablation study on the effectiveness of Negative Constraints across two backbones. HR denotes Hallucination Rate ($\downarrow$ lower is better); Perf. denotes downstream performance ($\uparrow$ higher is better).
Backbone Method M3CoT ScienceQA LLaVA-W
HR ($\%$) $\downarrow$ Acc. $\uparrow$ HR ($\%$) $\downarrow$ Acc. $\uparrow$ HR ($\%$) $\downarrow$ ROUGE-L $\uparrow$
Chameleon-7B w/o Constraints 5.8 30.1 4.5 52.5 8.4 27.9
AIM-CoT (Ours) 1.9 31.4 1.2 53.1 2.7 29.8
Qwen2-VL-7B w/o Constraints 3.5 43.9 2.8 56.9 5.1 35.1
AIM-CoT (Ours) 0.8 44.7 0.5 57.4 1.5 36.3
The results are summarized in Table 19, from which two key observations are derived:
First, the CAG module exhibits high intrinsic reliability across backbones. Even without explicit negative constraints, the hallucination rates remain relatively low (e.g., $6\%$ on M3CoT for both models). This suggests that the foundational strategy of using VLM-generated descriptions as context is robust and not inherently prone to generating misleading information, regardless of the underlying model architecture.
Second, negative constraints provide a critical layer of safety. Despite the solid baseline, the introduction of negative constraints consistently suppresses residual hallucinations across all datasets and backbones. Notably, for Chameleon-7B on M3CoT, the constraints reduce the hallucination rate by approximately 67% (from 5.8% to 1.9%). This demonstrates that while the models are generally capable, the constraints effectively filter out subtle noise and “over-interpretations.” By driving the hallucination rate down to a negligible level, we ensure that the subsequent AVP module operates on a virtually noise-free foundation, thereby maximizing overall reasoning performance.
L.2 Effectiveness of Safety Instructions in DAT Integration
In Section 4.4, we introduce a robust integration strategy within the DAT module, utilizing specific Safety Instructions (e.g., instructing the model to treat inserted regions as “supplementary references” and verify semantic consistency) to mitigate the risk of integrating irrelevant or hallucinated visual cues. To validate the efficacy of this textual safeguard, we conduct a stress test involving Irrelevant Visual Injection.
Experimental Setup. We simulate a worst-case scenario where the upstream selection process fails completely. Whenever the DAT triggers a visual insertion, instead of inserting the semantically relevant region selected by AVP, we deliberately inject a random, irrelevant image patch from a different dataset sample. We then measure the model’s resilience under this interference using two configurations: (1) Naive Injection (w/o Safety): The irrelevant patch is inserted directly into the reasoning chain without any accompanying safety prompts. (2) AIM-CoT Integration (w/ Safety): The irrelevant patch is inserted accompanied by our proposed safety instructions. We employ the same backbones (Chameleon-7B, Qwen2-VL-7B) and benchmarks (M3CoT, ScienceQA, LLaVA-W) as in previous sections.
Evaluation Metrics. We report the Performance under Noise and the Performance Drop ($\Delta$) relative to the clean, standard AIM-CoT performance (where accurate regions are inserted). A smaller magnitude of $\Delta$ indicates greater robustness against misleading visual information.
Table 20: Ablation study on Safety Instructions under irrelevant visual injection. $\Delta$ indicates the performance drop compared to the standard AIM-CoT ($\downarrow$ lower drop/magnitude is better).
Backbone Method M3CoT ScienceQA LLaVA-W
Acc. $\Delta$ Acc. $\Delta$ ROUGE-L $\Delta$
Chameleon-7B Naive Injection (w/o Safety) 24.5 -6.9 46.2 -6.9 22.1 -7.7
AIM-CoT (w/ Safety) 29.9 -1.5 51.8 -1.3 28.4 -1.4
Qwen2-VL-7B Naive Injection (w/o Safety) 38.2 -6.5 50.9 -6.5 30.5 -5.8
AIM-CoT (w/ Safety) 43.5 -1.2 56.1 -1.3 35.1 -1.2
The results in Table 20 provide compelling evidence for the necessity of safety instructions:
First, unprotected visual insertion is highly risky. When irrelevant patches are naively injected, the models suffer severe performance degradation (e.g., $\sim$7% drop on M3CoT for Chameleon-7B). This confirms that without guidance, VLMs tend to over-trust visual tokens, attempting to reason over noise and consequently hallucinating or diverging from the correct logic path.
Second, Safety Instructions act as an effective semantic firewall. By simply instructing the model to verify consistency, the performance drop is drastically reduced to a marginal level (e.g., only -1.5% on M3CoT). This demonstrates that the safety instruction successfully cues the model to exercise critical judgment: when the inserted visual evidence conflicts with the textual context (as in this random injection case), the model chooses to disregard the visual noise and rely on its internal knowledge, thereby maintaining robust performance even in the presence of upstream errors.
Appendix M Deployment of AIM-CoT
M.1 Analysis of the Complexity of AVP Module
@@
The AVP module’s primary purpose is to dynamically and intelligently select salient sub-regions of an image during the text generation process. This is achieved by calculating the “information gain” that each potential sub-region offers, thereby allowing the model to “zoom in” on relevant visual details and generate more informed and contextually aware text.
@@
This formula highlights that, thanks to KV caching, the complexity scales linearly with sequence length N—a substantial improvement over the standard quadratic dependence. While the cost is also linear with respect to the number of candidates $N_{C}$ and selections $K$, our framework maintains strong deployability. This is because the AVP module is highly efficient at extracting crucial visual information from the candidate set. Our empirical results demonstrate that AIM-CoT achieves exceptional performance (as shown in Table 1) even when these hyper-parameters are kept at low levels (e.g., $K=3,N_{C}=6$, which is the default setting). In conclusion, we can approximate the complexity of AVP as follows:
@@
This demonstrates that the combination of architectural optimizations (batching and KV caching) and the high extractive efficiency of AVP ensures the module’s practicality, making it efficient and readily deployable in practice. As a comparison, the attention-driven selection method within the baseline model ICoT is with a complexity of $O(logK\cdot\Delta N+L\cdot N\cdot H)$, which also scales linearly with sequence length N. Although this is acknowledged to be lower than the complexity of AVP shown in Equation 16, our empirical results detailed in Appendix M.2 suggest that AVP’s average inference time is no more than 1.36 times that of this method, while achieving performance far superior to it as analyzed in Section 5.4.
M.2 Empirical Analysis of the Deployment of AIM-CoT
@@
In this section, we empirically investigate the deployability of the AIM-CoT framework and the temporal overhead introduced by the AVP module. For the experimental setup, to ensure the reliability of the experimental results, we introduced two representative VLM backbones as the basis for the experiment: Chameleon-7B and Qwen2.5-VL-32B. They not only differ significantly in the number of parameters, but also employ different early-fusion and late-fusion strategies, respectively. In the 0-shot setting, we compare the average inference time of AIM-CoT against ICoT Gao et al. (2025). ICoT, as a key baseline model in this study, employs a Top-K strategy to simultaneously select regions with the highest attention scores for constructing the multimodal CoT, and is therefore expected to possess a relatively lower time complexity compared to AIM-CoT. Consequently, the comparison with ICoT serves as a direct indicator of AIM-CoT’s deployability.
Table 21: Comparison of average inference time (seconds). Values in parentheses denote the relative time against ICoT.
Dataset Method Chameleon-7B Qwen2.5-VL-32B
M3CoT ICoT 11.62 51.21
[…]
LLaVA-W ICoT 8.58 36.95
[…]
@@
The experimental results are presented in Table 21. Two key observations can be drawn. First, the AVP module does not introduce significant temporal costs to the AIM-CoT framework, an efficiency attributable to batch processing and the KV Cache mechanism. Second, AIM-CoT achieves substantially superior performance as shown in Table 1, at a time cost comparable to that of the efficient baseline, which is less than $1.36$ times that of ICoT. This suggests that our proposed AIM-CoT framework achieves a favorable trade-off between performance and deployability.